Find the exact value of tan 300 degrees. Use the sum or difference identity.
Found 2 solutions by mathmate, ikleyn:
Answer by mathmate(429)

Question:
Find the exact value of tan 300 degrees. Use the sum or difference identity.

Solution:
300 degrees = 5π/6
=π-π/6

and we know that tan(π)=0, and tan(π/6)=sqrt(3)/3
tan(300°)
=tan(π-π/6)
=(tan(π)-tan(π/6))/(1+tan(π)*tan(π/6))
=(0-sqrt(3)/3)/(1+0)
=-sqrt(3)/3

Answer by ikleyn(52787)

.
Question:
Find the exact value of tan 300 degrees. Use the sum or difference identity.

Solution:
300 degrees = 5π/3
=2π-(pi/3)

and we know that tan(2π)=tan(0)=0, and tan(π/3)=sqrt(3)
tan(300°)
=tan(2π-(pi/3))
=(tan(2π)-tan(π/3))/(1+tan(&2pi;)*tan(π/3))
=(0-sqrt(3))/(1+0)
=-sqrt(3)


The solution by other tutor is TOTALLY WRONG.